# Distributed quantum error correction based on hyperbolic Floquet codes

### v1

@@ Abstract
Quantum computing offers significant speedups, but the large number of physical qubits required for quantum error correction introduces engineering challenges for a monolithic architecture. One solution is to distribute the logical quantum computation across multiple small quantum computers, with non-local operations enabled via distributed Bell states. Previous investigations of distributed quantum error correction have largely focused on the surface code, which offers good error suppression but poor encoding rates, with each surface code instance only able to encode a single logical qubit. In this work, we argue that hyperbolic Floquet codes are particularly well-suited to distributed quantum error correction for two reasons. Firstly, their hyperbolic structure enables a high number of logical qubits to be stored efficiently. Secondly, the fact that all measurements are between pairs of qubits means that each measurement only requires a single Bell state. Through simulations, we show that distributed hyperbolic Floquet codes offer good performance under local and non-local phenomenological noise. This shows that distributed quantum error correction is not only possible but efficiently realisable.
@@
distributed quantum computing, quantum error correction
@@ I Introduction
To reach the millions of physical qubits required by industrially-relevant applications, a promising approach is a distributed architecture, where small quantum processing units are interconnected by quantum channels. By using these channels to share entanglement between processors, the quantum circuit required to implement the QEC code can be distributed across the network [18]. By using a distributed design, the number of qubits in the system can be scaled without increasing the complexity of the individual quantum processors. Furthermore, if these quantum channels are used to generate Bell states, then qubit teleportation or two-qubit gates can be performed using a shared Bell state, local operations, and classical communication [19].
Some approaches to distributed quantum computation consider a setup in which each quantum processor holds a separate surface code patch [20, 21]. In this approach, the network is only used for multi-qubit logical operations. An unwanted side-effect of this is that logical errors can occur along the edge of each quantum processing unit. This is problematic as errors are more likely to occur on communication qubits—physical qubits that are used to create the noisier non-local entanglement.
@@
In this work, we investigate the feasibility of implementing hyperbolic Floquet codes in a distributed setting. Starting with a hyperbolic Floquet code (part of which is shown in Fig. 1a), we use the fine-graining approach from [29] to increase the distance of the code (Fig. 1b). Finally, we distribute this code across multiple quantum processors of a fixed size, as shown in Fig. 1c. Through simulations, we show that these codes are fault-tolerant under experimentally feasible numbers for both local and non-local phenomenological noise, and derive resource requirements for a quantum memory experiment of equivalent length to a million logical operations (“MegaQuOp”) [31, 32]. This paves the way for implementing efficient quantum error-correcting codes in a distributed setting.
@@ III-A Hyperbolic and semi-hyperbolic codes
Consider a graph with $n$ vertices (or data qubits) and genus $g$. The number of logical qubits that can be encoded is denoted by $k=2g$ and the encoding rate of a code is denoted by $k/n=2g/n$. As mentioned earlier, the hyperbolic codes considered have efficient encoding rates but only achieve logarithmic scaling for code distance $d$. For example, the $\{8,3\}$ hyperbolic lattice has genus $g=n/16+1$. Thus, the encoding rate is $k/n=1/8+2/n$ which is constant in the limit $n\rightarrow\infty$, whereas the distance $d$ grows as $\log(n)$[30]. Using the fine-graining procedure in [29], the rate-distance trade-off can be adjusted in order to improve the distance at the expense of adding more qubits. Fine-graining a hyperbolic code adds planar patches to its surface, and hence the generated codes families are termed semi-hyperbolic [29, 17]. Since fine-graining preserves the topological properties of the lattice, the encoding rate is unchanged, while the distance can be increased to at most $O(\sqrt{n})$. Fig. 3 shows the fine-graining process in a small area of an $\{8,3\}$ octagonal tiling. To perform fine-graining, we begin with the dual of the lattice. In the dual lattice, the vertices are located on the faces of the initial lattice, and the edges are drawn between them. As a result, qubits that are on vertices in the original lattice are represented by triangles in the dual. Fine-graining involves further triangulating each of these triangles, i.e., each triangle in the dual graph is replaced with a triangle-bounded interior-triangular lattice of size $f$. The level of fine-graining can be increased by increasing $f$. The semi-hyperbolic lattice is obtained by taking the dual of this fine-grained lattice. In the case of $\{8,3\}$ tiling, the semi-hyperbolic lattice consists of both hexagons and octagons. Regarding the nomenclature, we represent hyperbolic codes as “Hx” where x is the number of data qubits. Semi-hyperbolic codes are named “Hx-fy”, where Hx indicates the base hyperbolic code with x data qubits, and f signifies fine-graining level y.
@@ III-B Quantum code partitioning for modular embedding
In a physical implementation, we would use a cut-off wait time for the Bell state. In a Floquet code, a QEC round comprises three sub-rounds, namely $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$ checks. If the wait time is shorter than the QEC sub-round, we will carry out further checks. If the wait time exceeds the QEC sub-round, we will mark it as an error and continue with additional checks.
@@ IV Simulation results
We numerically estimated the logical error rates for the family of hyperbolic and semi-hyperbolic Floquet codes described in Section III. The circuits were constructed and simulated using Stim, a tool for simulating quantum error correction circuits [51]. Errors detected by the Floquet codes considered here can be decomposed into weight-two “graph-like” errors, meaning that decoding can be performed using Minimum Weight Perfect Matching (MWPM) [52]. We used a phenomenological noise model which also accounts for the challenges of running a distributed quantum code across a network, details of which are provided in Section IV-A.
From these simulations we estimated pseudo-thresholds for a phenomenological noise model. Furthermore, we investigated the effect of non-uniform noise introduced by partitioning the code across multiple devices. Finally, we estimate the number of physical qubits required to provide a MegaQuOp quantum memory for our distributed hyperbolic Floquet codes.
@@ IV-A Noise model
The development of quantum error correction aims to identify a specific process of operations to protect quantum information stored in a noisy environment. Quantum memory experiments are performed based on an error model to reflect potential real errors that may occur, ensuring that theoretical successes can be translated into practical applications. In this work, we assume that the primary source of error stems from an uncontrolled environment or through the quantum network, rather than imprecise qubit gate tuning. Hence, the error models are defined in terms of depolarising channels, rather than Pauli error biases. We employ a modified phenomenological noise model as our primary error model.
@@
In a distributed setting, non-local weight-two checks are performed using a shared heralded Bell state that serves as an auxiliary qubit(s) for error detection [53]. It is reasonable to assume that the fidelities of these Bell states during initialization will not be comparable with the initialization fidelities of the single auxiliary qubits within the quantum processing unit. Additionally, we must consider the time delay associated with generating the Bell state. Below, we outline a method to incorporate these noise parameters into the simulations.
@@
Under the standard phenomenological noise model controlled by a single error parameter $p_{\textrm{local}}$, each single-qubit (resp. two-qubit) operation is followed by a single-qubit (resp. two-qubit) depolarising error with probability $p_{\textrm{local}}$. Each weight-two measurement is implemented directly using the MPP gate in Stim [51], and followed by a two-qubit depolarising channel with probability $p_{\textrm{local}}$. There is also a chance of a measurement result being incorrectly recorded, which happens with probability $p_{\textrm{local}}$.
In addition to standard phenomenological noise assumptions, we also consider noisy Bell state generation for non-local checks. Non-local checks are modelled using the MPP gate in Stim with a elevated measurement flip probability of $p_{\textrm{non-local}}$.
@@
In each of the $X\otimes X$, $Y\otimes Y$, or $Z\otimes Z$ QEC sub-rounds, the weight-two checks are distributed across the network according to the partitioning discussed in Section III-B. We then wait for the formation of Bell states corresponding to non-local checks to complete the QEC sub-round. Generating heralded non-local entanglement is inherently stochastic, necessitating multiple attempts to successfully generate a Bell state [31]. The wait time $\tau$ to produce a Bell state is measured in gate cycles, with one gate cycle defined as the duration needed to execute a single qubit gate. We suppose that $\tau=30$ gate cycles, which is roughly the time required to perform a single QEC sub-round [54]. We assume that the data qubits idle during the wait time and experience decoherence. Between each QEC sub-round, each data qubit idles and is considered to undergo $\tau$ layers of single-qubit depolarising noise. For example, during a complete cycle of $X\otimes X$, $Y\otimes Y$ and $Z\otimes Z$ checks, $\tau$ layers of depolarising noise are applied to every data qubit after each QEC sub-round. We perform simulations by repeating detector rounds, where a detector round comprises 2 QEC rounds (6 sub-rounds).
@@
Table I provides our full noise model in terms of the gates available in Stim [51]. Here MPP stands for two-qubit Pauli measurements of the form $P\otimes P$. We assume negligible cross-talk or correlated errors, i.e., applying a gate on qubit 1 does not change the errors experienced by qubit 2. We also assume that we are working in the bulk of the computation, meaning that preparation and measurement of the logical qubits are assumed to be perfect.
@@
TABLE I: Error probabilities of gate types under the phenomenological noise model
@@
Single-qubit gate $p_{\textrm{local}}$
@@
Measurement (intra-QPU) MPP ($p_{\textrm{local}}$)
Measurement (inter-QPU) MPP ($p_{\textrm{non-local}}$)
Wait time (per QEC sub-round) $\tau$ = 30
Decoherence (per QEC sub-round) $1-(1-p_{\textrm{local}})^{\tau}$
@@ IV-B Pseudo-threshold
Figure 4: Pseudo-threshold for H64-f3. Logical error rates are shown for a single worst-case observable in the code H64-f3 (576 data qubits) partitioned over 32 qubit devices for 64 detector rounds. The error bars are roughly the same size as the symbols.
@@
The estimation process starts by performing a 512-point grid scan of the 2-D region defined by $p_{\textrm{local}}\in 10^{[-4.5,-2.5]}$ and $p_{\textrm{non-local}}\in 10^{[-3,-1]}$. We sample until either $256$ errors or $2\times 10^{5}$ total shots have been taken for each point. We fit a proxy Gaussian process to this data using [55]. We use individual proxy models for the logical error rates of both the H64-f3 code and a physical qubit. We derive a difference model, defined as $|\log(p_{\textrm{encoded}})-\log(p_{\textrm{unencoded}})|$, where $p_{\textrm{encoded}}$ and $p_{\textrm{unencoded}}$ denote the probabilities of a logical error for a logical qubit encoded in the H64-f3 Floquet code and a single physical qubit, respectively. This function of each logical error probability will be zero at the pseudo-threshold and allows us to use robust search processes to find physical error rates on the pseudo-threshold.
@@
We perform 3 steps of refinement sampling using local penalization with a batch size of 8 to refine the estimated pseudo-threshold. We perform this sampling within fixed "contexts", corresponding to points drawn from 2 linear scans over the $p_{\textrm{local}}$ and $p_{\textrm{non-local}}$ ranges. This ensures we find many minima, corresponding to points that cover the entire pseudo-threshold curve. The 2-D logical error rate landscape was interpolated onto a $800\times 800$ point regular grid with SciPy’s griddata function [56] and smoothed using a Gaussian kernel with size 7. Initial threshold estimates were drawn from this smoothed landscape (plotted prior to smoothing as background colour in Fig. 4.
@@
Starting from each initial estimate, we perform narrow sweeps to identify the pseudo-threshold with higher precision. These sweeps are performed until we have either $10^{5}$ errors or $5\times 10^{7}$ shots. We sweep over the recorded pseudo-threshold in both the $p_{\textrm{local}}$ and $p_{\textrm{non-local}}$ axes for each point. We then perform a linear fit to both the physical and H64-f3 data from those sweeps, noting that the data is approximately linear due to the narrow sweep range. We take the point of intersection of those fits as the threshold. The error in the fits is shown as error bars on the threshold points.
@@
The pseudo-threshold is shown in Fig. 4. Red points indicate fitted locations of the pseudo-threshold, and the background shading corresponds to the error ratio between H64-f3 and a physical qubit for a quantum memory experiment. In the distributed setting, a Floquet encoded qubit can tolerate a higher $p_{\textrm{non-local}}$ error compared to $p_{\textrm{local}}$. The intuitive reason for the higher noise tolerance of the non-local connections is that they represent a small fraction of the entanglement resources used in performing the parity checks. Hence, when an error is detected, the decoder is more likely to associate it with the non-local link, which enhances the tolerance for non-local errors.
@@ IV-C Sub-threshold performance
where $\epsilon_{L}$ is the logical error rate per detector round where a detector round comprises 2 QEC rounds (6 sub-rounds), $d$ is the code distance, and $C$ is a fitting constant. $\epsilon_{L}$ is determined by expressing the overall logical error rate over a quantum memory experiment as
@@
Figure 5: Figures show (left) the logical error rate for code families with error rates fixed at $p_{\textrm{local}}=99.99\%$ and $p_{\textrm{non-local}}=99.5\%$. The sub-threshold error correction scaling $\Lambda$ (right) is calculated using Equation 1. The data points in the right plot are derived from the error rate fits in the left plot.
[…]

### v2

@@ Abstract
Quantum computing offers significant speedups, but the large number of physical qubits required for quantum error correction introduces engineering challenges for a monolithic architecture. One solution is to distribute the logical quantum computation across multiple small quantum computers, with non-local operations enabled via distributed Bell states. Previous investigations of distributed quantum error correction have largely focused on the surface code, which offers good error suppression but poor encoding rates, with each surface code instance only able to encode a single logical qubit. In this work, we argue that hyperbolic Floquet codes are particularly well-suited to distributed quantum error correction for two reasons. Firstly, their hyperbolic structure enables a high number of logical qubits to be stored efficiently. Secondly, the fact that all measurements are between pairs of qubits means that each measurement only requires a single Bell state. Through simulations, we show that distributed hyperbolic Floquet codes offer good performance under local and non-local circuit-level noise. This shows that distributed quantum error correction is not only possible but efficiently realisable.
@@
quantum networking, distributed quantum computing, quantum error correction
@@ I Introduction
To reach the millions of physical qubits required by industrially-relevant applications, a promising approach is a distributed architecture, where fixed size quantum processing units are interconnected by quantum channels. By using these channels to share entanglement between processors, the quantum circuit required to implement the QEC code can be distributed across the network [18]. By using a distributed design, the number of qubits in the system can be scaled without increasing the complexity of the individual quantum processors. Furthermore, if these quantum channels are used to generate Bell states, then qubit teleportation or two-qubit gates can be performed using a shared Bell state, local operations, and classical communication [19].
Some approaches to distributed quantum computation consider a setup in which each quantum processor holds a separate surface code patch [20, 21, 22]. In this approach, the network is only used for multi-qubit logical operations. An unwanted side-effect of this is that logical errors can occur along the edge of each quantum processing unit. This is problematic as errors are more likely to occur on communication qubits—physical qubits that are used to create the noisier non-local entanglement.
@@
In this work, we investigate the feasibility of implementing hyperbolic Floquet codes in a distributed setting. Starting with a hyperbolic Floquet code (part of which is shown in Fig. 1a), we use the fine-graining approach from [30] to increase the distance of the code (Fig. 1b). Finally, we distribute this code across multiple quantum processors of a fixed size, as shown in Fig. 1c. Through simulations, we show that these codes are fault-tolerant under experimentally feasible numbers for both local and non-local circuit noise, and derive resource requirements for a quantum memory experiment of equivalent length to a million logical operations (“MegaQuOp”) [32, 33]. This paves the way for implementing efficient quantum error-correcting codes in a distributed setting.
@@ III-A Hyperbolic and semi-hyperbolic codes
Consider a graph with $n$ vertices (or data qubits) and genus $g$. The number of logical qubits that can be encoded is denoted by $k=2g$ and the encoding rate of a code is denoted by $k/n=2g/n$. As mentioned earlier, the hyperbolic codes considered have efficient encoding rates but only achieve logarithmic scaling for code distance $d$. For example, the $\{8,3\}$ hyperbolic lattice has genus $g=n/16+1$. Thus, the encoding rate is $k/n=1/8+2/n$ which is constant in the limit $n\rightarrow\infty$, whereas the distance $d$ grows as $\log(n)$[31]. Using the fine-graining procedure in [30], the rate-distance trade-off can be adjusted in order to improve the distance at the expense of adding more qubits. Fine-graining a hyperbolic code adds planar patches to its surface, and hence the generated codes families are termed semi-hyperbolic [30, 17]. Since fine-graining preserves the topological properties of the lattice, the number of encoded logical qubits remains unchanged, while the distance can be increased by at most $O(\sqrt{n})$. Fig. 3 shows the fine-graining process in a small area of an $\{8,3\}$ octagonal tiling. To perform fine-graining, we begin with the dual of the lattice. In the dual lattice, the vertices are located on the faces of the initial lattice, and the edges are drawn between them. As a result, qubits that are on vertices in the original lattice are represented by triangles in the dual. Fine-graining involves further triangulating each of these triangles, i.e., each triangle in the dual graph is replaced with a triangle-bounded interior-triangular lattice of size $f$. The level of fine-graining can be increased by increasing $f$. The semi-hyperbolic lattice is obtained by taking the dual of this fine-grained lattice. In the case of $\{8,3\}$ tiling, the semi-hyperbolic lattice consists of both hexagons and octagons. Regarding the nomenclature, we represent hyperbolic codes as “Hx” where x is the number of data qubits. Semi-hyperbolic codes are named “Hx-fy”, where Hx indicates the base hyperbolic code with x data qubits, and fy signifies fine-graining level y.
@@ III-B Quantum code partitioning for modular embedding
In a physical implementation, we would use a cut-off wait time for the Bell state. In a Floquet code, a plaquette forming round comprises three sub-rounds, namely $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$ checks. If the wait time is shorter than the sub-round, we will carry out further checks. If the wait time exceeds the sub-round, we will mark it as an error and continue with additional checks.
@@ IV Simulation results
We numerically estimated the logical error rates for the family of hyperbolic and semi-hyperbolic Floquet codes described in Section III. The circuits were constructed and simulated using Stim, a tool for simulating quantum error correction circuits [52]. Errors detected by the Floquet codes considered here can be decomposed into weight-two “graph-like” errors, meaning that decoding can be performed using Minimum Weight Perfect Matching (MWPM) [53]. We used a circuit noise model which also accounts for the challenges of running a distributed quantum code across a network, details of which are provided in Section IV-A.
From these simulations we estimated pseudo-thresholds for a circuit noise model. Furthermore, we investigated the effect of non-uniform noise introduced by partitioning the code across multiple devices. Finally, we estimate the number of physical qubits required to provide a MegaQuOp quantum memory for our distributed hyperbolic Floquet codes.
@@ IV-A Noise model
The development of quantum error correction aims to identify a specific process of operations to protect quantum information stored in a noisy environment. Quantum memory experiments are performed based on an error model to reflect potential real errors that may occur, ensuring that theoretical successes can be translated into practical applications. In this work, we assume that the primary source of error stems from an uncontrolled environment or through the quantum network, rather than imprecise qubit gate tuning. Hence, the error models are defined in terms of depolarising channels, rather than Pauli error biases. We employ a modified circuit-level noise model as our primary error model.
@@
In a distributed setting, non-local weight-two checks are performed using a shared heralded Bell state that serves as an auxiliary qubit(s) for error detection [54]. It is reasonable to assume that the fidelities of these Bell states during initialization will not be comparable with the initialization fidelities of the single auxiliary qubits within the quantum processing unit. Additionally, we must consider the time delay associated with generating the Bell state. In the following, we outline a method to incorporate these noise parameters into the simulations.
@@
In the standard circuit-level noise model controlled by a single error parameter $p_{\textrm{local}}$, each single-qubit operation (state preparation, gate, and reset) is followed by a single-qubit depolarising channel with error probability $p_{\textrm{local}}$. Each two-qubit gate is followed by a two-qubit depolarising channel with error probability $p_{\textrm{local}}$. We apply a depolarising channel with the $p_{\textrm{local}}$ error probability prior to a qubit measurement. There is a chance that a measurement result is incorrectly recorded, which occurs with probability $p_{\textrm{local}}$. In addition to standard assumptions, we also consider noisy Bell state generation for non-local checks. We assume that a depolarising noise with probability $p_{\textrm{non-local}}$ is applied after the creation of the Bell state.
@@
In each of the $X\otimes X$, $Y\otimes Y$, or $Z\otimes Z$ plaquette forming rounds, the weight-two checks are distributed across the network according to the partitioning discussed in Section III-B. We then wait for the formation of Bell states corresponding to non-local checks to complete the plaquette forming round. Generating heralded non-local entanglement is inherently stochastic, necessitating multiple attempts to successfully generate a Bell state [32]. We suppose that number of attempts required to initialise a Bell state is 5 gate cycles, as a consequence of a assumed 20% success rate of the cavity enhanced and heralded remote bell pair generation scheme[55]. We assume that the data qubits remain idle while waiting to be involved in an operation during which they experience decoherence. During this time, they encounter $t$ layers of depolarizing noise with probability $p_{\textrm{local}}$, where $t$ represents the waiting time for each qubit which can be kept track. We perform simulations by repeating detector rounds, where a detector round comprises 2 plaquette forming rounds (6 sub-rounds).
@@
Table I provides our full noise model in terms of the gates available in Stim [52]. We assume negligible cross-talk or correlated errors, i.e., applying a gate on qubit 1 does not change the errors experienced by qubit 2.
@@
TABLE I: Error probabilities of gate types under the circuit noise model
@@
Single-qubit (prepare, gate, reset, measure) $p_{\textrm{local}}$
@@
Remote Bell state prepare $p_{\textrm{non-local}}$
Decoherence (per circuit step) $p_{\textrm{local}}$
@@ IV-B Pseudo-threshold
Figure 4: Pseudo-threshold for H144-f3. Logical error rates are shown for a single worst-case observable in the code H144-f3 (1296 data qubits) partitioned over 32 qubit devices for 12 detector rounds (72 colour rounds). Error bars in the threshold fitting are shown as solid red lines around threshold points. The ratio of logical to physical error rate is shown as a 2D color map.
@@
The estimation process starts by performing a 512-point grid scan of the 2-D region defined by $p_{\textrm{local}}\in 10^{[-4,-2.5]}$ and $p_{\textrm{non-local}}\in 10^{[-3,-0.5]}$. We sample until either $512$ errors or $2\times 10^{5}$ total shots have been taken for each point. We fit a proxy Gaussian process to this data using [56]. We use individual proxy models for the logical error rates of both the H144-f3 code and a physical qubit. We derive a difference model, defined as $|\log(p_{\textrm{encoded}})-\log(p_{\textrm{unencoded}})|$, where $p_{\textrm{encoded}}$ and $p_{\textrm{unencoded}}$ denote the probabilities of a logical error for a logical qubit encoded in the H144-f3 Floquet code and a single physical qubit, respectively. This function of each logical error probability will be zero at the pseudo-threshold and allows us to use robust search processes to find physical error rates on the pseudo-threshold.
@@
We perform 7 steps of refinement sampling using local penalization with a batch size of 10 to refine the estimated pseudo-threshold. We perform this sampling within fixed "contexts", corresponding to points drawn from 2 linear scans over the $p_{\textrm{local}}$ and $p_{\textrm{non-local}}$ ranges. This ensures we find many minima, corresponding to points that cover the entire pseudo-threshold curve. The 2-D logical error rate landscape was interpolated onto a $800\times 800$ point regular grid with SciPy’s griddata function [57] and smoothed using a Gaussian kernel with size 7. Initial threshold estimates were drawn from this smoothed landscape. The interpolated landscape is plotted as the 2D colour map in Fig. 4.
@@
Starting from each initial estimate, we perform narrow sweeps to identify the pseudo-threshold with higher precision. These sweeps are performed until we have either $100$ errors or $4\times 10^{7}$ shots. We sweep over the recorded pseudo-threshold in both the $p_{\textrm{local}}$ and $p_{\textrm{non-local}}$ axes for each point. We then perform a linear fit to both the physical and H144-f3 data from those sweeps, noting that the data is approximately linear due to the narrow sweep range. We take the point of intersection of those fits as the threshold. The error in the fits is shown as error bars on the threshold points.
@@
The pseudo-threshold is shown in Fig. 4. Red points indicate fitted locations of the pseudo-threshold, and the 2D color map corresponds to the error ratio between H144-f3 and a physical qubit for a quantum memory experiment. In the distributed setting, a Floquet encoded qubit can tolerate a higher $p_{\textrm{non-local}}$ error compared to $p_{\textrm{local}}$. At the highest tested local error rate, we find a pseudothreshold at $96.9\pm 3.7\%$ remote bell state preparation fidelity and $99.99\pm 0.02\%$ local fidelity. The pseudothreshold curve traces a smooth landscape through to $99.71\pm 0.56\%$ remote bell state fidelity and a $99.95\pm 0.14\%$ local fidelity. The intuitive reason for the higher noise tolerance of the non-local connections is that they represent a small fraction of the entanglement resources used in performing the parity checks. Hence, when an error is detected, the decoder is more likely to associate it with the non-local link, which enhances the tolerance for non-local errors.
@@ IV-C Sub-threshold performance
where $\epsilon_{L}$ is the logical error rate per detector round where a detector round comprises 2 plaquette forming rounds (6 sub-rounds), $d$ is the code distance, and $C$ is a fitting constant. $\epsilon_{L}$ is determined by expressing the overall logical error rate over a quantum memory experiment as
@@
Figure 5: Figures show (left) the logical error rate for code families with error rates fixed at $p_{\textrm{local}}=99.97\%$ and $p_{\textrm{non-local}}=99\%$. The sub-threshold error correction scaling $\Lambda$ (right) is calculated using Equation 1. The data points in the right plot are derived from the error rate fits in the left plot.
For the error model under evaluation, $\Lambda=5.06\pm 0.88$ indicates the potential for a strong exponential suppression of logical error rates through modular scaling. H400 provides 52 logical qubits, and so in order to perform an identity MegaQuOp, that is, $10^{6}$ logical operations each requiring a detector round with a cumulative any-observable 50% chance of failure, a system would require an observable-level success rate greater than $\exp(\log(0.5)/104)=99.3\%$ after $1923010^{6}/52=$ 19230 detector rounds. This implies a failure rate of $3.5\times 10^{-5}$ per detector round. This failure rate is shown on Fig. 5 as a black horizontal line. At the tested error rates, H400-f3 would satisfy this requirement using 5400 qubits including measurement ancilla, in under 170 32-qubit devices equipped with 8 qubit-photon interfaces per device.